A rectangle´s length is 8 cm more than three times its width. If the perimeter is 128 cm, find
the length and width.

Answer by checkley77(12844)

L=3W+8
2L+2W=128
2(3W+8)+2W=128
6W+16+2W=128
8W=128-16
8W=112
W=112/8
W=14 ANS. FOR THE WIDTH.
L=3*14+8
L=42+8
L=50 ANS. FOR THE LENGTH.
PROOF:
2*50+2*14=128
100+28=128
128=128